# The Agony of Opacity: Foundations for Reflective Interpretability in AI-Mediated Mental Health Support

### v1

@@ 5 Tensions and Needs
An evaluation of level of reflection (and its impacts) is particularly important in the context of adding new features or functionality. In the case of a reflective interpretability approach, it is possible that adding additional information for a distressed user to reflect on might actually disincentivize engaging with AI-mediated mental health support. Multiple studies [song2024typing, siddals2024happened, jung2025ve] describe how a major incentive for engaging with AI-mediated mental health support was the ease at which users felt able to find support from general purpose chatbots, and the level of agency they felt with regards to how that support progressed (including when and where they could access support). Adding additional scaffolding, such as prompting a user to input information about their close friends (as part of an advance directive) or guiding them through support boundaries, could make users feel overwhelmed or surveilled, and result in an additional barrier to accessing support. More qualitative research is needed, including interviews with service users, to better understand what forms of scaffolding would not make accessing support less acceptable to users in distress. However, as an intermediate step, interface designers could use measures of agency and intervention acceptability from past mental health research (such as the State Hope Scale [snyder1996development] or the Program Feedback Scale [schleider2019virtual]) to evaluate whether the addition of new reflective interpretability mechanisms served the purpose of increasing a user’s feelings of agency in their care (the State Hope Scale) without reducing their willingness to engage with support (the Program Feedback Scale).
@@
At a broader level, this research could feed into policy guidelines around AI-mediated mental health support that apply across general purpose chatbot companies. In clinical mental health practice, clinical regulations require clinicians to clearly define boundaries around the type of support that can be provided (as part of the informed consent process) [fisher2008informed, trachsel2019strengthen], and ensure that users are well-aware of how their information or data might be used [pendse2024advancing]. More research work is needed to understand where including that information would actually enrich the process of accessing support for users in distress, and at what level that might become so burdensome that it disincentivizes care.
We therefore call for lightweight and opt-in reflective mechanisms, with efficacy measured via established scales and augmented through qualitative work, towards balancing interpretability against the practical need to ensure support is accessible for those who have few other options.

### v2

@@ 5 Tensions and Potential Risks
Further empirical research is necessary to better understand what kinds of design approaches to presenting this information to users spur a greater sense of reflection and interpretation. An evaluation of level of reflection (and its impacts) is thus particularly important in the context of adding new features or functionality. More qualitative research is needed, including interviews with service users, to better understand what forms of scaffolding would not make accessing support less acceptable to users in distress. However, as an intermediate step, interface designers could use measures of agency and intervention acceptability from past mental health research (such as the State Hope Scale [95] or the Program Feedback Scale [84]) to evaluate whether the addition of new reflective interpretability mechanisms serves the purpose of increasing a user’s feelings of agency in their care (the State Hope Scale) without reducing their willingness to engage with support (the Program Feedback Scale). In addition, it is important to also consider how the process of interpretation and reflection can be extremely diverse across cultures, including when associated with technology use. The creation of culturally-sensitive scales to measure reflection and interpretation of outputs, drawn from research on cultural differences in how people interpret new information [19] and news [91], could be helpful alongside qualitative interviews. A slower and more reflective use of chatbots for AI-mediated mental health support might also be less costly (through reduced usage), and metrics around total engagement and usage could be incorporated into measures of reflective interpretability.
@@
At a broader level, this research could feed into policy guidelines around AI-mediated mental health support that apply across general purpose chatbot companies. In clinical mental health practice, clinical regulations require clinicians to clearly define boundaries around the type of support that can be provided (as part of the informed consent process) [31, 108], and ensure that users are well-aware of how their information or data might be used [77]. More research work is needed to understand where including that information would actually enrich the process of accessing support for users in distress, and at what level that might become so burdensome that it disincentivizes care. For instance, recent regulations have required AI chatbots to disclose to users that they are speaking to an AI chatbot (and that the user is not speaking to a human) every three hours [58]—research around reflective interpretability could help inform similar regulations, towards allowing users to feel more informed (and a greater sense of agency) when interacting with AI-mediated mental health support.
@@ 6 Conclusion
Appendix A: Key Insights for Reflective Interpretability
In this appendix, we provide core takeaways and key insights from this paper.
Question Key Takeaway
What is reflective interpretability? We understand reflective interpretability to be an agency-preserving and iterative process of reflecting on model outputs related to one’s experience of distress, and interpreting that information towards creating meaning from it. We delineate three core aspects of reflective interpretability, occurring throughout engagement with AI-mediated mental health support: 1) the ability of users to actively make sense of how model outputs were generated, 2) clarity on the fixed boundaries of support that can be provided by a model (including what kinds of interactions might trigger those boundaries), and 3) functionality that encourages users to reflect on model outputs and interpret outputs in ways that promote long-term well-being
Why is it important to take into consideration in AI-mediated mental health support? Past research has demonstrated how the acute experience of distress can alter an individual’s perspective in ways that make opacity particularly dangerous. A state of acute distress may make it easier to accept new information as fact, and act on it as directive advice, particularly if it appears to be coming from an authority figure (as is often the case with AI chatbot interfaces).
What are the conceptual precedents of reflective interpretability? The conceptual precedents of reflective interpretability originate in research from medical ethics around informed consent being a reflective and interpretative dialogue between provider and client, centered on active discussion of needs and boundaries, alongside research from HCI around interface design that centers user agency and authority (rather than that of the technical system).
Why do we call it reflective interpretability? The practice of mental health support fundamentally resists providing objective explanations of behavior or experience to individuals, and encourages client reflection and interpretation, with the client acting as an active expert in their healing process.
How might reflective interpretability be integrated into the interfaces associated with AI-mediated mental health support? Role induction (from psychotherapy), prosocial advance directives (from crisis intervention), intervention titration (from psychiatry), and mechanisms for recourse (from care authorization).
What potential tensions or harms exist alongside a reflective interpretability approach? And how might these be mitigated? Adding more reflective interpretability features to an interface might cause additional friction in an environment where friction might disincentivize engaging with (potentially life-saving) support. In addition, showing too much of the model’s internal reasoning could be perceived as manipulative or offensive. More research is needed to understand how best to integrate these features such that they do not make getting support more difficult.